# New flexible versions of extended generalized Pareto distribution for count data

### v2

@@ 1 Introduction
Many statistical models are available for analyzing non-negative count data. Standard frameworks include the Poisson distribution, which suits count data with equal mean and variance, and the Negative Binomial distribution, which addresses overdispersion by allowing variance to exceed the mean. For data with an excessive number of zeros, ZI models such as the ZI Poisson (ZIP) and ZI Negative Binomial (ZINB) extend these distributions to account for additional zero inflation. The DGPD, on the other hand, is tailored for modeling discrete exceedances above a high threshold, making it ideal for analyzing extreme values and rare events. Each model addresses different data characteristics and analysis needs, providing versatile tools for statistical analysis. In practice, however, count data are often overdispersed to these distributions, and sometimes extreme observations are also present. For instance, the count data of upheld complaints of automobile insurance companies complaints rankings in New York City and the number of doctors visit to the hospital are the heavier tail datasets and have excessive zeros. In both datasets, 47.53% and 42% are zero values, meaning no upheld complaints registered and no doctors visit is registered in the hospital, and 4.63% and 4.91% are extreme observations that occur above the 95% quantile of the datasets, respectively. We fitted ZINB to both datasets and constructed their quantile-quantile (Q-Q) plots. Figure 1 (a, b) shows that the model fits well to the lower and central parts of the data distribution but exhibits deficiencies at the upper tail. Consequently, building a model for the count data sets with ZI and extreme observations becomes a significant problem and has not yet been properly studied.
@@
Figure 1: Example of a two-part figure with individual sub-captions showing that captions are flush left and justified if greater than one line of text.
Modeling discrete exceedances is a viable approach to analyzing extreme tail behavior, offering valuable insights into rare events. The peak-over-threshold (POT) method, as introduced by Pickands, (1975), commonly approximates the distribution of exceedances above a high threshold to Generalized Pareto Distribution (GPD). Several adaptations of the GPD for discrete data are available in the literature (Krishna and Pundir,, 2009, Buddana and Kozubowski,, 2014, Kozubowski et al.,, 2015). More recently, Hitz et al., (2024) discussed the DGPD for modeling the tail behavior of integer-valued random variables. Consequently, when the threshold is sufficiently high, standard extreme value models, such as the POT method with DGPD approximation (Ranjbar et al.,, 2022, Hitz et al.,, 2024), can be applied to model the exceedances. Detailed information is provided in Section 2. Choosing an appropriate threshold $u$ is critical for effectively applying the POT approximation. Setting the threshold too low can introduce bias into the estimates, as the DGPD is justified only in an asymptotic sense. On the other hand, setting the threshold too high reduces the number of data points, increasing the estimation variance. In practice, selecting a suitable threshold in a continuous setting often involves parameters stability plots and mean residual life plots, as suggested by Coles, (2001, ch. 4) and Davison and Smith, (1990). A suitable threshold can also be identified where the mean of the exceedances (mean residual life) starts to exhibit a linear relationship with the threshold value (Daouia et al.,, 2023). In the discrete POT approach, defining a threshold at a high quantile remains necessary, which poses challenges due to the discrete nature of the data.
@@ 2 Modeling framework
$\Pr(Y\leq y)=\mathcal{G}\left(\mathcal{F}\left({y+1};\beta,\xi\right)\right),$ (7)
@@
The distribution defined by (9) is referred to as the ZIDEGPD. The explicit formula of CDF of ZIDEGPD is written as
$\Pr(Y\leq y)=\pi+(1-\pi)\mathcal{G}\left(\mathcal{F}\left(m+1,\beta,\xi\right)\right),\qquad y\in\mathbb{N}_{0},$ (10)
@@
$f_{u}(u;\kappa)=\left\{\begin{array}[]{ll}\frac{2\sqrt{\kappa}}{2\Phi\left(\sqrt{\kappa}\right)-1}\phi\left(\sqrt{\kappa}(u-1)\right)&\text{if}\hskip 14.22636pt\kappa>0\\ 1&\text{if}\hskip 14.22636pt\kappa\rightarrow 0.\end{array}\right.$
The rest of the article will refer to the model associated with this $\mathcal{G}(.)$ choice as DEGPD-Normal and ZIDEGPD-Normal.
@@
We consider truncated Beta distribution as another choice to add more flexibility in the density of $U$. Let $D_{x}(\alpha_{1},\alpha_{2})$ indicates incomplete beta function:
@@
and the density function of Beta distribution is written as
@@
We choose $\omega=1/32$ here. Interestingly, this particular choice yields sufficient flexibility for the tail of the distribution. On the other hand, if the selection of $\omega$ for some event is not flexible, we can choose another value in the interval $(0,1/2)$ or could be estimated from the observed data. Nevertheless, this may cause identifiability problems in the later scenario. This model will be referred to as DEGPD-Beta and ZIDEGPD-Beta in the paper.
@@ 5.2 ZIEGPD
We assess the performance of the proposed ZIDEGPD models ($M_{1}$, $M_{2}$, and $M_{3}$) fitted to the DV dataset. The results of these fitted models are presented in Table 2. One notable observation is that the proportion of zeros, represented by the parameter $\pi$, is accurately estimated across all models. The estimated values of $\pi$ are very close to the true proportion of zeros in the dataset, demonstrating the ZIDEGPD model’s ability to correctly capture the zero inflation in the data. There is variation across the models regarding the bulk part parameter $\kappa$. This variability arises because each model separately estimates the effect of zero values through the parameter $\pi$, which somehow influences the distribution of the bulk part of the data. The tail index parameter, $\xi$, remains relatively consistent across all models. This suggests that despite differences in bulk part estimation, the tail behavior of the data, captured by $\xi$, is similarly modeled in all cases. Furthermore, the BIC value of $M_{3}$ indicates it performs slightly better than $M_{1}$ and $M_{3}$. Figure 8 (top) presents the Q-Q plots for the ZIDEGPD fitted models, illustrating the best fit of models ($M_{1}$, $M_{2}$, and $M_{3}$) to the zero-inflated heavy tail data.

### v3

@@ 1 Introduction
Many statistical models are available for analyzing non-negative count data. Standard frameworks include the Poisson distribution, which suits count data with equal mean and variance, and the Negative Binomial distribution, which addresses overdispersion by allowing variance to exceed the mean. For data with an excessive number of zeros, ZI models such as the ZI Poisson (ZIP) and ZI Negative Binomial (ZINB) extend these distributions to account for additional zero inflation. The DGPD, on the other hand, is tailored for modeling discrete exceedances above a high threshold, making it ideal for analyzing extreme values and rare events. Each model addresses different data characteristics and analysis needs, providing versatile tools for statistical analysis. In practice, however, count data are often overdispersed to these distributions, and sometimes extreme observations are also present. For instance, the count data of upheld complaints of automobile insurance companies complaints rankings in New York City and the number of doctors visits to the hospital are heavy-tailed datasets and have excessive zeros. In both datasets, 47.53% and 42% of the values are zeros, indicating no upheld complaints registered and no recorded doctor visits at the hospital, respectively. Conversely, 6.18% and 2.48% of the observations are extreme values in both datasets, respectively, identified using the right fence with $c=3$ of the Tukey fence method (Schwertman et al.,, 2004). We fitted ZINB to both datasets and constructed their quantile-quantile (Q-Q) plots. Figure 1 (a, b) shows that the model fits well to the lower and central parts of the data distribution but exhibits deficiencies at the upper tail. Consequently, building a model for the count data sets with ZI and extreme observations becomes a significant problem and has not yet been properly studied.
@@
Figure 1: The Q-Q plots of fitted ZINB to the count data sets. Upheld complaints of the automobile insurance companies rankings of New York City (left), and doctors visits to the hospital (right).
Modeling discrete exceedances is a viable approach to analyzing extreme tail behavior, offering valuable insights into rare events. The peak-over-threshold (POT) method, as introduced by Pickands, (1975), commonly approximates the distribution of exceedances above a high threshold to the Generalized Pareto Distribution (GPD). Several adaptations of the GPD for discrete data are available in the literature (Krishna and Pundir,, 2009, Buddana and Kozubowski,, 2014, Kozubowski et al.,, 2015). More recently, Hitz et al., (2024) discussed the DGPD for modeling the tail behavior of integer-valued random variables. Consequently, when the threshold is sufficiently high, standard extreme value models, such as the POT method with DGPD approximation (Ranjbar et al.,, 2022, Hitz et al.,, 2024), can be applied to model the exceedances. Detailed information is provided in Section 2. Choosing an appropriate threshold $u$ is critical for effectively applying the POT approximation. Setting the threshold too low can introduce bias into the estimates, as the DGPD is justified only in an asymptotic sense. On the other hand, setting the threshold too high reduces the number of data points, increasing the estimation variance. In practice, selecting a suitable threshold in a continuous setting often involves parameter stability plots and mean residual life plots, as suggested by Coles, (2001, ch. 4) and Davison and Smith, (1990). A suitable threshold can also be identified where the mean of the exceedances (mean residual life) starts to exhibit a linear relationship with the threshold value (Daouia et al.,, 2023). In the discrete POT approach, defining a threshold at a high quantile remains necessary, which poses challenges due to the discrete nature of the data.
@@ 2 Modeling framework
$\Pr(Y\leq y)=\mathcal{G}\left(\mathcal{F}\left({y+1};\beta,\xi\right),\kappa\right),$ (7)
@@
The distribution defined by (9) is referred to as the ZIDEGPD. The explicit formula of the CDF of ZIDEGPD is written as
$\Pr(Y\leq y)=\pi+(1-\pi)\mathcal{G}\left(\mathcal{F}\left(y+1,\beta,\xi\right),\kappa\right),\qquad y\in\mathbb{N}_{0},$ (10)
@@
$f_{u}(u;\kappa)=\left\{\begin{array}[]{ll}\frac{2\sqrt{\kappa}}{2\Phi\left(\sqrt{\kappa}\right)-1}\phi\left(\sqrt{\kappa}(u-1)\right)&\text{if}\hskip 14.22636pt\kappa>0\\ 1&\text{if}\hskip 14.22636pt\kappa\rightarrow 0,\end{array}\right.$
where $\phi$ is the density function of the standard normal distribution. The rest of the article will refer to the model associated with this $\mathcal{G}(.)$ choice as DEGPD-Normal and ZIDEGPD-Normal.
@@
We consider the truncated Beta distribution as another choice to add more flexibility in the density of $U$. Let $D_{x}(\alpha_{1},\alpha_{2})$ indicates incomplete beta function:
@@
and the density function of the Beta distribution is written as
@@
Figure 2: Flexibility check for model $M_{3}$ with varying $\omega$ values and keeping the values of other parameters as $\kappa=5,\beta=1,$ and $\xi=0.1$. In left panel $\omega\in(0,1/16)$ while in right panel $\omega\in(1/8,1/2.01)$, where $2.01$ ensure the choice of $\omega$ is not exactly 0.5.
To choose the appropriate value of $\omega$, we have check the flexibility of the model $M_{3}$ in Figure 2 with different values of $\omega$. In the left panel, it can be seen that the model is flexible when varying the values of $\omega$ from $(0,1/64,1/48,1/32,1/16)$. The behavior is almost similar in most of the distribution in the bulk part, and the tail is constant. On the other hand, when varying $\omega$ $(1/8,1/2)$, the model loses flexibility in the bulk part of the distribution. It turns out that $\omega=1/32$ leads to a flexible model for low values with no change in the right tail. If the model is not flexible enough in some real applications, the value could be set to another value, as shown in the right panel of Figure 2, or could be estimated from the data. Nevertheless, this may cause identifiability problems in the later scenario. This model will be referred to as DEGPD-Beta and ZIDEGPD-Beta in the paper.
@@ 5.2 ZIEGPD
We assess the performance of the proposed ZIDEGPD models ($M_{1}$, $M_{2}$, and $M_{3}$) fitted to the DV dataset. The results of these fitted models are presented in Table 2. One notable observation is that the proportion of zeros, represented by the parameter $\pi$, is accurately estimated across all models. The estimated values of $\pi$ are very close to the true proportion of zeros in the dataset, demonstrating the ZIDEGPD model’s ability to correctly capture the zero inflation in the data. There is variation across the models regarding the bulk part parameter $\kappa$. This variability arises because each model separately estimates the effect of zero values through the parameter $\pi$, which somehow influences the distribution of the bulk part of the data. The tail index parameter, $\xi$, remains relatively consistent across all models. This suggests that despite differences in bulk part estimation, the tail behavior of the data, captured by $\xi$, is similarly modeled in all cases. Furthermore, the BIC value of $M_{3}$ indicates it performs slightly better than $M_{1}$ and $M_{2}$. Figure 9 (top) presents the Q-Q plots for the ZIDEGPD fitted models, illustrating the best fit of models ($M_{1}$, $M_{2}$, and $M_{3}$) to the zero-inflated heavy tail data.